# MSVD-Indonesian: A Benchmark for Multimodal Video-Text Tasks in Indonesian

### v1

@@ Abstract
Multimodal learning on video and text data has been receiving growing attention from many researchers in various research tasks, including text-to-video retrieval, video-to-text retrieval, and video captioning. Although many algorithms have been proposed for those challenging tasks, most of them are developed on English language datasets. Despite Indonesian being one of the most spoken languages in the world, the research progress on the multimodal video-text with Indonesian sentences is still under-explored, likely due to the absence of the public benchmark dataset. To address this issue, we construct the first public Indonesian video-text dataset by translating English sentences from the MSVD dataset to Indonesian sentences. Using our dataset, we then train neural network models which were developed for the English video-text dataset on three tasks, i.e., text-to-video retrieval, video-to-text retrieval, and video captioning. The recent neural network-based approaches to video-text tasks often utilized a feature extractor that is primarily pretrained on an English vision-language dataset. Since the availability of the pretraining resources with Indonesian sentences is relatively limited, the applicability of those approaches to our dataset is still questionable. To overcome the lack of pretraining resources, we apply cross-lingual transfer learning by utilizing the feature extractors pretrained on the English dataset, and we then fine-tune the models on our Indonesian dataset. Our experimental results show that this approach can help to improve the performance for the three tasks on all metrics. Finally, we discuss potential future works using our dataset, inspiring further research in the Indonesian multimodal video-text tasks. We believe that our dataset and our experimental results could provide valuable contributions to the community. Our dataset is available on GitHub [^1] https://github.com/willyfh/msvd-indonesian.
Keywords video-text dataset  $\cdot$ video captioning  $\cdot$ video-to-text retrieval  $\cdot$ text-to-video retrieval  $\cdot$ neural network
@@ 1 Introduction
Multimodal machine learning [1] aims to construct machine learning models that have the ability to learn from multiple modalities, such as natural language, visual, and audio. The recent attention on multimodal machine learning has led to new research problems connecting video and text modalities, such as text-to-video retrieval [2], video-to-text retrieval [3], and video captioning [4]. While the text-to-video retrieval task aims to retrieve a relevant video given a text query as an input, the video-to-text retrieval task aims to retrieve a relevant text given a video query as an input. For the video captioning task, the objective is to generate a natural language sentence describing the input video. In order to develop a machine learning model for the multimodal video-text tasks, a video-text dataset consisting pair of videos and texts is needed.
The majority of video-text datasets [5, 6, 7] are constructed with English sentences. Only a few of the datasets are publicly available in a non-English language, e.g., Chinese [8], Turkish [9], Italian [10], and Hindi [11]. Despite the Indonesian language is one of the most widely used languages in the world, to the best of our knowledge, there is no public video-text dataset available for the language at the time of writing this paper. Consequently, the research progress of the multimodal video-text for the Indonesian language is still under-explored. In response, we construct the first public Indonesian video-text dataset, by translating English sentences from Microsoft Video Description (MSVD) dataset [5] to Indonesian sentences. Our MSVD-Indonesian dataset consists of 1970 videos with roughly 80k sentences in total, which is the same as the MSVD dataset. Using our novel dataset, we then build neural network models which were developed and evaluated on the English video-text dataset on three video-text tasks, i.e., text-to-video retrieval, video-to-text retrieval, and video captioning. Our dataset and the experimental results may well be groundbreaking for the research of multimodal video-text in the Indonesian language on the three tasks.
Employing neural networks algorithms along with the pretrained feature extractors is one of the popular approaches in video-text tasks [12, 13, 14, 15]. While it is possible that the same approach on the English dataset may also be effective on the Indonesian dataset, the actual performance of the algorithm when applied to Indonesian languages is still unknown. Especially when the model or the feature extractor was mainly pretrained on an English language dataset, the applicability of the algorithm on the Indonesian language is even more questionable due to the lack of pretraining visual-language resources, i.e., pretrained model and pretraining dataset. In our experiments, we consider a scenario when the pretraining resources of visual-language for the Indonesian language are not available. To overcome the unavailability of the pretraining resources, we apply transfer learning by utilizing the models pretrained on the English vision-language dataset as the feature extractors, and then train the overall models on our MSVD-Indonesian dataset. By working on this scenario, we provide insight on how is the performance of language-specific video-text algorithms which were developed for English, when it is trained on a language which does not have the pretraining resources. Note that incorporating a pretraining process using an Indonesian dataset is beyond the scope of this study, which could be explored in future research. In addition, we discuss some potential future works in utilizing our dataset in section 6.
In this paper, to investigate the performance of video-text algorithms on our dataset, we employ X-CLIP [15], which is one of the state-of-the-art models in text-to-video retrieval and video-to-text retrieval tasks on the MSVD dataset. For the video captioning task, we employ VNS-GRU [12], which is one of the state-of-the-art models on the MSVD dataset. Following X-CLIP and VNS-GRU, we utilize the visual-language feature extractors, i.e., contrastive language-image pre-Training (CLIP) [16] and semantic concept detection (SCD) model [17], which was pretrained on English datasets. We then fine-tune the X-CLIP and the VNS-GRU models and perform several ablation studies to investigate the effect of some configurations of the algorithms on our MSVD-Indonesian dataset.
@@
  • We construct and release a video-text dataset in the Indonesian language, by translating an existing English video-text dataset, i.e., MSVD. To the best of our knowledge, this is the first publicly available video-text dataset in the Indonesian language at the time of writing this paper.
  • We establish the baseline performance of our dataset for three video-text tasks, i.e., text-to-video retrieval, video-to-text retrieval, and video captioning, by training neural network-based methods on the MSVD-Indonesian dataset.
  • Our experimental results show that the cross-lingual pretraining and fine-tuning can help to improve the performance in the three video-text tasks for the Indonesian language on all metrics.
  • Finally, we present some possible future works using our dataset to inspire other researchers in leading to a broader understanding of the topic and new discoveries.
2 Related Works
@@ 2.1 Video-Text Datasets
The video-text datasets are often constructed to facilitate research and development in multimodal learning, specifically for tasks like video captioning [4] and video-text retrieval [15]. The sentences in the video-text datasets are commonly annotated in the English language, with only a few available in the non-English language [8, 9, 10, 11]. The MSVD dataset [5], which is one of the most popular video-text datasets, was mainly available in the English language, with few subsequent variants constructed in non-English languages, i.e., MSVD-CN [18] and MSVD-Turkish [9]. The original MSVD dataset initially came in multiple languages, which includes, English, Hindi Romanian, Slovene, Serbian, Tamil, Dutch, German, etc. However, since the dataset was not publicly available anymore, Chen and Dolan [5] recollected the dataset, and able to reconstruct only the English sentences. Later on, the Chinese version [18] and the Turkish version [9] of video-text datasets are built based on the reconstructed MSVD dataset. In this paper, we construct and release a video-text dataset with Indonesian sentences based on the English MSVD dataset. Making our dataset publicly available encourages innovation by providing a resource for researchers to develop new ideas in the area of multimodal video-text in the Indonesian language.
@@ 2.2 Video-Text Retrieval
Video-text retrieval consists of two main problems, i.e., text-to-video and video-to-text retrieval. Given a text as an input query, a text-to-video retrieval model retrieves the relevant video as the output. On the other hand, a video-to-text retrieval model receives a video as the input query, then retrieves the relevant text as the output. While the goals of the two tasks are different, many algorithms developed are often able to be used for both retrieving videos and retrieving text given the text query and the video query, respectively. Luo et al. [13] introduced a method called CLIP4Clip, by utilizing a CLIP model [16] which was pretrained on a large-scale image-language dataset. They showed that the usage of the CLIP model is able to achieve a SOTA performance for both text-to-video and video-to-text retrieval tasks with a significant margin. Ma et al. [15] then extends the CLIP4Clip model by proposing a multi-grained contrastive model, namely X-CLIP, which helped to even further improve the performance. Although those works have shown that utilizing the pretrained CLIP model could boost the accuracy performance on the video-text datasets in the English language, its applicability to the Indonesian language is still unknown due to the lack of the pertaining resources for the Indonesian language. In this work, we adopt X-CLIP and apply transfer learning to transfer the knowledge of the pretrained English CLIP model by fine-tuning the model to our Indonesian video-text dataset. Through our experiments, we obtain performance results for both text-to-video retrieval and video-to-text retrieval on our dataset, providing insights into how the pretrained English CLIP model can be useful for these two tasks in the Indonesian language. In addition, our experimental results can be served as a benchmark for future research in Indonesian video-text retrieval tasks.
@@ 2.3 Video Captioning
Video captioning aims to generate a sentence given a video as input. Many recent methods in video captioning utilized feature extractors pretrained on vision and language datasets, e.g., CLIP [16] and SCD [17]. While the CLIP model is pretrained on pairs of images and sentences, the SCD model is pretrained on pairs of videos and keywords in the sentences. Gan et al. [17] introduced SCD for video captioning to capture semantic features from the videos. They pretrained SCD in multi-label settings, where the video is used as the input and the keywords in the sentences are used as the labels. The pretrained SCD is then used as the semantic feature extractor to extract the probabilities of the keywords given the video as the input. Borrowing the success of SCD, some other works [12, 19, 20, 21] adopted the SCD-based approach as the feature extractor in their proposed methods. Chen et al. [19] proposed a network called semantic detection network (SDN), adopting the idea of SCD, i.e., pretrained on video and keywords in the sentences. In another work, Chen et al. [12] incorporate a similar network called tagging network, which is also basically utilizing the concept of SCD. Similarly, Perez Martin et al. [20, 21] incorporated SCD into a customized LSTM-based decoder, which additionally receives the semantic features as the input. In this paper, we adopt the VNS-GRU algorithm [12], which has the best performance among those SCD-based algorithms for video captioning on the MSVD dataset. We first pretrain the SCD-based feature extractor using the pair of video and English keywords, then we train the overall algorithm on the Indonesian MSVD dataset. The performance results of this experiment establish the benchmark performance for future studies of video captioning in the Indonesian language.
@@ 3.1 Dataset Collection
The initial version of the MSVD dataset consists of 2089 videos [5]. However, due to some videos being removed from YouTube, Chen et al. were able to recollect and archive only 1970 out of 2089 videos. Later on, this reconstructed MSVD dataset, i.e., the one with 1970 videos, is the version that is popularly used for video-text research tasks in the English language. In this work, we collect the 1970 videos along with the 80827 English sentences from the MSVD dataset, and then translate each of the sentences using Google Translate API to have the Indonesian sentences. Thus, for each video in our dataset, we have the same number of sentences as in the MSVD dataset. Because every sentence in our dataset has a one-to-one mapping with the sentence in the English MSVD dataset, our dataset can also be used for the research of multilingual video-text research.
@@ 3.2 Dataset Analysis
Using a machine translation service to translate sentences may lead to inaccuracies in the translations due to the limitations of the service. Since we utilize Google Translate API, we found that the Indonesian sentences in our dataset could have inaccurate grammar. Although some of the translations are inaccurate, many of the sentences for each video are still well-translated. As shown in Figure 1 (a), we can see that the three samples of the ground truth sentences in the MSVD-Indonesian dataset, i.e., indicated by ID-#, are translated well w.r.t. to the English ground truth, i.e., indicated by EN-#. In Figure 1 (b), however, we can notice that the verb ‘flours’ in EN-2 is inaccurately translated as ‘tepung’ in ID-2. The machine translation is unable to translate the verb ‘flours‘ correctly since it does not have a direct equivalent verb in Indonesian. The correct translation for the sentence could be ‘seorang wanita menuangkan tepung ke dalam wajan (a woman pours flour into the pan)‘. The inaccuracy of the sentence in our dataset is not solely due to the problem of the machine translation service. In a few cases, the English sentences from the MSVD dataset originally consist of inaccurate annotations. From the same Figure 1 (b), we can notice that there is an inaccurate annotation in the English annotation, i.e., EN-3. Despite there being no strawberry in the video, the video is inaccurately annotated with ‘make strobery food’. Interestingly, although there is a typographical error in the English annotation, i.e., ‘strobery’, the machine translation translates the inaccurate English word ‘strobery’ as an accurate Indonesian word ‘stroberi (Strawberry)’. In our experiments, we retain the inaccurate sentences, including both inaccurate translations and annotations, without making any modifications. We consider the low-quality annotations as noise and outliers of our dataset.
Figure 1: Two samples of videos from the MSVD-Indonesian dataset, along with their corresponding annotated sentences. Each video is accompanied by three sample sentences.
As illustrated in Figure 2, we further analyze and compare the sentences in the MSVD dataset and the MSVD-Indonesian dataset on several data visualizations. In Figure 2 (a), we can see similar patterns on the most frequent words in both datasets. For instance, the articles ‘seorang’, ‘seekor’, and ‘sebuah’ are frequently used in the MSVD-Indonesian dataset, which is consistent with the MSVD dataset where similar articles are also frequently used, i.e., ’a’ and ’an’. The different frequencies for the articles between the two datasets are due to the usage of the articles being different. While the choice between ’a’ and ’an’ depends on the sound that follows, i.e., consonant or vowel, the choice among ‘seorang’, ‘seekor’, and ‘sebuah’ depends on the specific category of the noun that follows them, i.e., human, animal, or general item. From Figure 2 (b), we can notice that the MSVD dataset has more vocabulary words compared to the MSVD-Indonesian dataset. We found that the MSVD dataset has 12592 unique words, while the MSVD-Indonesian dataset has 9457 unique words. In terms of the average sentence length, as shown in Figure 2 (c), the MSVD-Indonesian dataset has about 5.7 words and the MSVD dataset has about 7 words. These statistical differences may indicate that a video-text algorithm that performs the best for the MSVD dataset, is not automatically the best for our MSVD-Indonesian dataset.
Figure 2: Comparison of MSVD and MSVD-Indonesian dataset using three different visualizations: (a) word cloud, (b) bar chart of vocabulary word count, and (c) bar chart of average sentence length.
@@ 4 Methods
The methods that we adopt in this work are the ones that utilized a feature extractor pretrained on the English vision-language dataset. For the text-to-video and video-to-text retrieval tasks, we utilized X-CLIP [15] which incorporates the CLIP [16] model pretrained on an image-text dataset. As for the video captioning task, we employ VNS-GRU [12] which utilizes the SCD [17] model pretrained on a video-text dataset. Assuming there are no available pretraining resources for the vision-language in Indonesian, we initially use the feature extractors, i.e., CLIP and SCD, which were primarily pretrained on the English datasets. Subsequently, using the English-based pretrained extractors, we train the X-CLIP and VNS-GRU models on the MSVD-Indonesian dataset for their respective tasks.
@@ 4.1 X-CLIP
As the extension of CLIP4Clip [13], X-CLIP [15] also utilized a pretrained CLIP model which was mainly pretrained on the English language. However, different than CLIP4Clip which only considers video-sentence contrastive learning, X-CLIP employed multi-grained contrastive learning where four different types of contrastive are considered, i.e., video-sentence, video-word, sentence-frame, and frame-word. To train the X-CLIP model, we need to extract frame-level features, video-level features, word-level features, and sentence-level features using the pretrained CLIP model. Following the same procedure as in X-CLIP, we first sample the video frames with a sampling rate of 1 FPS. Given the sampled frames, we then extract the frame-level feature using the pretrained CLIP model by inputting the video frames into the visual encoder of the CLIP model. Different than X-CLIP, we do not capture the interaction between frames using the temporal encoder [15] since our experiments show that the temporal encoder hurts the performance of the model in our dataset. The video-level features are then computed by averaging the frame-level features in the video. For extracting word-level and sentence-level features, the text encoder of the CLIP model is used by taking the outputs of the tokens from the final layer of the text encoder, where the [EOS] token output is used as the sentence-level features. All of these extracted features are then used to compute the similarity score of the four different types of contrastive relations in the X-CLIP algorithm. Especially for relations of video-word, sentence-frame, and frame-word, the attention over similarity matrix (AOSM) module [15] is applied to fuse the similarity vector or matrix, thereby obtaining the instance-level similarity. Please refer to [15] for more details explanations of the X-CLIP model.
@@ 4.2 VNS-GRU
VNS-GRU [12] is one of the methods which adopts a pretrained SCD model [17], i.e., tagging network in [12], as a semantic feature extractor. The SCD model is a model which is pretrained under multi-label settings, where the training data are pairs of a video and keywords. The keywords are constructed by considering the top $n$ vocabulary from the English sentences in the video-text dataset. Given a video as the input, the pretrained SCD model is able to extract a semantic feature vector that represents the probability of keywords for the given video input. In the VNS-GRU approach, this semantic feature is then concatenated with another feature, which is extracted from the final layer of a multi-class classification model pretrained on ImageNet [22]. The concatenation of these features is then used as the final semantic features. Along with these semantic features, VNS-GRU also requires video-level features which are extracted from a 3D CNN model pretrained on an action recognition dataset, i.e., Kinetics-400 [23]. Using the extracted features, the VNS-GRU model is then trained by adopting three techniques [12], i.e., incorporating variational dropout and layer normalization in RNN unit, comprehensive selection method, and professional learning. For a more comprehensive understanding of the VNS-GRU model, please refer to [12].
@@ 5.1.2 Captioning
For the captioning experiments, we assess the performance of the models by using four popular metrics in video captioning, i.e., BLEU@4 [24]. ROUGE-L [25], METEOR [26], and CIDEr [27]. 1) BLEU@4 computes the accuracy of a method by taking the precision of the generated sentences in terms of 4-grams, i.e., sequence of 4 words. 2) ROUGE-L measures the harmonic mean of precision and recall on the longest common subsequence (LCS) between generated sentence and ground-truth sentence. 3) METEOR computes its score by utilizing a weighted F-score based on unigrams, and incorporating a penalty function to penalize the incorrect word order in the generated sentence. 3) CIDER utilizes a voting-based method to have a robust measurement against noise or incorrect annotations.

### v2

@@ Abstract
Multimodal learning on video and text has seen significant progress, particularly in tasks like text-to-video retrieval, video-to-text retrieval, and video captioning. However, most existing methods and datasets focus exclusively on English. Despite Indonesian being one of the most widely spoken languages, multimodal research in Indonesian remains under-explored, largely due to the lack of benchmark datasets. To address this gap, we introduce the first public Indonesian video-text dataset by translating the English captions in the MSVD dataset into Indonesian. Using this dataset, we evaluate neural network models which were developed for the English video-text dataset on three tasks, i.e., text-to-video retrieval, video-to-text retrieval, and video captioning. Most existing models rely on feature extractors pretrained on English vision-language datasets, raising concerns about their applicability to Indonesian, given the scarcity of large-scale pretraining resources in the language. We apply a cross-lingual transfer learning approach by leveraging English-pretrained extractors and fine-tuning models on our Indonesian dataset. Experimental results demonstrate that this strategy improves performance across all tasks and metrics. We release our dataset publicly to support future research and hope it will inspire further progress in Indonesian multimodal learning [^1] https://github.com/willyfh/msvd-indonesian.
@@ 1 Introduction
Multimodal machine learning [1] enables models to learn from multiple modalities such as text, vision, and audio. Recent advances in this field have led to progress in video-text tasks, including text-to-video retrieval [23], video-to-text retrieval [16], and video captioning [25]. These tasks typically rely on supervised training over large-scale datasets consisting of paired video and textual descriptions.
However, the majority of available video-text dataset, e.g., MSVD [4], MSR-VTT [28], and ActivityNet Captions [10], are constructed in English. Only a few multilingual datasets exist for languages such as Chinese [26], Turkish [30], Italian [20], and Hindi [21]. Despite Indonesian being one of the most spoken languages globally, there is currently no publicly available video-text dataset for it. This lack of resources limits the development and evaluation of multimodal systems for Indonesian.
To address this gap, we construct the first public Indonesian video-text dataset by translating English captions in the MSVD dataset into Indonesian. Our MSVD-Indonesian dataset consists of 1,970 videos and approximately 80,000 Indonesian sentences, mirroring the structure and scale of the original dataset.
We use our dataset to evaluate neural network models originally developed for English video-text tasks. Specifically, we adopt X-CLIP [13] for retrieval and VNS-GRU [5] for captioning. These models rely on pretrained feature extractors like CLIP [19] or semantic concept detection (SCD) [7], which are trained on predominantly English data. To adapt them for Indonesian, we apply cross-lingual transfer learning by reusing English-pretrained extractors and fine-tuning them on our dataset.
Our experiments demonstrate that this transfer learning approach improves performance across all three tasks. We release the MSVD-Indonesian dataset publicly and hope it encourages further research in multilingual and low-resource video-text learning.
@@
  • We release the first Indonesian video-text dataset, translated from MSVD.
  • We establish baseline results for three tasks using models originally developed for English video-text tasks.
  • We demonstrate that cross-lingual transfer learning is effective for the Indonesian video-text tasks.
  • We outline future directions enabled by this dataset.
2 Related Work
@@ 2.1 Video-Text Datasets
Video-text datasets support a variety of multimodal tasks, such as video captioning [25] and retrieval [13]. Most existing datasets, including MSVD [4], MSR-VTT [28], and ActivityNet Captions [10], are constructed with English annotations. A few multilingual datasets exist, including MSVD-CN [14], MSVD-Turkish [30], and others in Chinese, Hindi, and Italian [26, 20, 21]. Although the original MSVD dataset was collected in multiple languages, the publicly released version by Chen and Dolan [4] includes only English. Later efforts reconstructed Chinese and Turkish versions from the English set. To our knowledge, no Indonesian version exists. Our work addresses this by constructing and releasing MSVD-Indonesian, the first public video-text dataset in Indonesian.
@@ 2.2 Video-Text Retrieval
Video-text retrieval encompasses text-to-video and video-to-text retrieval tasks. Recent methods have benefited significantly from vision-language pretraining. Luo et al. proposed CLIP4Clip [22], which uses CLIP [19] to encode video frames and text into a shared space. Ma et al. extended this with X-CLIP [13], introducing multi-granular contrastive learning to improve retrieval performance. While these methods perform well on English benchmarks, their cross-lingual effectiveness remains untested. In our work, we apply X-CLIP to our Indonesian dataset using cross-lingual transfer learning and assess its effectiveness on both retrieval tasks.
@@ 2.3 Video Captioning
In video captioning, many models use pretrained semantic extractors. One such model is the Semantic Concept Detector (SCD) by Gan et al. [7], which predicts keywords from videos in a multi-label fashion. SCD-based approaches have been incorporated into various captioning models [5, 6, 17, 18]. Among them, VNS-GRU [5] achieves strong results on MSVD. In this work, we adopt VNS-GRU and adapt it to our Indonesian dataset using SCD pretrained on English annotations. This setup allows us to investigate whether semantic features learned in English can transfer to a low-resource language scenario.
@@ 3.1 Dataset Collection
The MSVD dataset originally contained 2089 videos [4], but due to removed YouTube links, only 1970 videos were retained and widely adopted in prior research. We use this reconstructed version, including 80,827 English sentences, and translate all sentences to Indonesian using the Google Translate API. Each video in our dataset has the same number of sentences as in MSVD, preserving a one-to-one mapping. This alignment enables multilingual research and cross-lingual benchmarking.
@@ 3.2 Dataset Analysis
Machine-translated data can contain grammatical or semantic inaccuracies. While many Indonesian sentences are translated well, others contain artifacts. Figure 1(a) shows samples where the translations closely match the English captions. Figure 1(b) shows issues such as mistranslating the verb “flours” to “tepung.” In some cases, errors originate from the original English annotations, such as a mislabeled mention of ”strobery” that doesn’t appear in the video. Interestingly, the translation of “strobery” is corrected to the standard Indonesian “stroberi.” We retain such noisy samples to reflect real-world dataset imperfections.
Figure 1: Samples from the MSVD-Indonesian dataset. Each video is shown with three English (EN-#) and Indonesian (ID-#) sentence pairs.
We also analyze vocabulary size and sentence length. As shown in Figure 2, the MSVD-Indonesian dataset contains 9,457 unique words, compared to 12,592 in English. The average sentence length is 5.7 words in Indonesian and 7 words in English. These differences suggest that models trained on MSVD-English may not directly transfer well to Indonesian, highlighting the need for adaptation and re-evaluation on our dataset.
Figure 2: Comparison of MSVD and MSVD-Indonesian datasets based on (a) vocabulary size and (b) average sentence length.
@@ 4 Methods
We adopt two models previously applied to English video-text tasks that rely on pretrained vision-language features: X-CLIP [13] for text-video retrieval, and VNS-GRU [5] for video captioning. Our setup assumes a low-resource scenario in which large-scale vision-language pretraining data for Indonesian is unavailable. Accordingly, we use CLIP [19] and SCD [7] pretrained on English data, and fine-tune both models on our MSVD-Indonesian dataset.
@@ 4.1 X-CLIP
X-CLIP extends CLIP4Clip [22] by introducing multi-granular contrastive learning across video-sentence, video-word, sentence-frame, and frame-word levels. We follow its standard setup and extract frame features from 1 FPS sampled frames using the CLIP visual encoder. Sentence- and word-level features are obtained from the text encoder. Unlike the original model, we omit the temporal encoder, as our experiments show it degrades performance on our dataset. Final video features are computed by averaging frame embeddings. Cross-modal similarities are fused using the attention-over-similarity module (AOSM) [13]. We refer readers to the original paper for architectural details.
@@ 4.2 VNS-GRU
VNS-GRU integrates semantic features from an SCD model [7], which is pretrained to predict frequent keywords in a video. The output semantic vector is concatenated with features from an ImageNet-pretrained classifier. Additionally, video features are extracted using a 3D CNN pretrained on Kinetics-400 [8]. The final GRU-based decoder is trained using techniques such as variational dropout, layer normalization, and comprehensive selection, following [5]. We retain the original architecture and focus on evaluating its effectiveness in the Indonesian language setting.
@@ 5.1.2 Captioning
For the captioning experiments, we assess the performance of the models by using four popular metrics in video captioning, i.e., BLEU@4 [15]. ROUGE-L [11], METEOR [2], and CIDEr [24]. 1) BLEU@4 computes the accuracy of a method by taking the precision of the generated sentences in terms of 4-grams, i.e., sequence of 4 words. 2) ROUGE-L measures the harmonic mean of precision and recall on the longest common subsequence (LCS) between generated sentence and ground-truth sentence. 3) METEOR computes its score by utilizing a weighted F-score based on unigrams, and incorporating a penalty function to penalize the incorrect word order in the generated sentence. 4) CIDER utilizes a voting-based method to have a robust measurement against noise or incorrect annotations.